At 100 K and 0.1 atmospheric pressure the volume of helium gas is 10 litres If the volume and pressure are doubled its temperature will change to

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

Use

 P1V1/T1 = P2 V2/T2

P1V1/T1   = 2P1 *2V1 /T2

T2=4T1

    =4 * 100

   =400 K

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin